OpenStudy (anonymous):

Find the slope and y-intercept of the line. for a medal y = - 4/3x - 3

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So what number is in the m position ? This is your slope. And what number is in the b position ? This is your y intercept

OpenStudy (idealist10):

y=mx+b where m=slope and b=y-intercept

OpenStudy (idealist10):

In your case, m=-4/3 and b=-3. So the slope is -4/3 and the y-intercept is -3.

OpenStudy (texaschic101):

exactly :)